Suppose demand for inkjet printers is estimated to be QX = 1000 – 5PX + 10PY – 2PZ + 0.1M. If own price (PX) = 80, related prices, PY = 50, PZ = 150, and income, M = 20,000.Calculate own price elasticity of demand and interpret your result


1
Expert's answer
2022-01-02T18:19:47-0500

Solution:

Own price elasticity of demand = "\frac{\triangle QX}{\triangle PX} \times \frac{PX}{QX}"

First, derive QX:

QX = 1000 – 5(80) + 10(50) – 2(150) + 0.1(20,000) = 1000 – 400 + 500 – 300 + 2,000 = 1,800

QX = 1,800


"\frac{\triangle QX}{\triangle PX}" = -5


PX = 80

Own price elasticity of demand = -5 "\times" "\frac{80}{1,800}"= -0.22

Own price elasticity of demand = -0.22


Own price elasticity of demand is less than 1, which means that the demand is price inelastic (demand is not sensitive to price changes).